### v1

# Towards a Unified Framework for Evaluating Explanations

Abstract
The challenge of creating interpretable models has been taken up by two main research communities: ML researchers primarily focused on lower-level explainability methods that suit the needs of engineers, and HCI researchers who have more heavily emphasized user-centered approaches often based on participatory design methods. This paper reviews how these communities have evaluated interpretability, identifying overlaps and semantic misalignments. We propose moving towards a unified framework of evaluation criteria and lay the groundwork for such a framework by articulating the relationships between existing criteria. We argue that explanations serve as mediators between models and stakeholders, whether for intrinsically interpretable models or opaque black-box models analyzed via post-hoc techniques. We further argue that useful explanations require both faithfulness and intelligibility. Explanation plausibility is a prerequisite for intelligibility, while stability is a prerequisite for explanation faithfulness. We illustrate these criteria, as well as specific evaluation methods, using examples from an ongoing study of an interpretable neural network for predicting a particular learner behavior.
keywords:
Explainable AI, evaluating explanations, model transparency, interpretable neural networks
1 Introduction
The growing awareness in educational data mining (EDM) of a need for more explainable AI (XAI) has led to the increasing adoption of interpretability methods. Such methods are continually being developed and refined within research communities such as machine learning (ML) and human-computer interaction (HCI). However, the critical task of evaluating the efficacy of the explanations created has not been sufficiently explored. Tellingly, a systematic review of explainable student performance prediction models did not find a single study that evaluated the explanations it produced [1]. Furthermore, a standardized framework for conducting such evaluations is still lacking [15].
In this position paper, we aim to foster discussion to begin addressing this gap by proposing the goal of a unified framework for evaluating explanations. We review concepts critical to the goals of XAI, including the intended contexts of an explanation, the multiple research milieux of explainability and intelligibility, and proposed evaluation methods. We argue that the evaluation of explanations should be based on a set of criteria that must be met for an explanation to be useful and propose a hierarchy of criteria that brings together some previously described in the literature. Finally, we illustrate these criteria using an ongoing study of an interpretable neural network for predicting a particular learner behavior. We conclude by discussing the implications of this initial perspective on an evaluation framework for future research in XAI.
2 What to evaluate?
The explainability literature has often highlighted the difference between intrinsically interpretable models that are designed with transparency in mind and opaque black-box models that require post-hoc explainability methods [12]. At face value, models and explanations seem like two very different objects to evaluate. However, even intrinsically interpretable models—such as linear regression models or decision trees—require some form of explanation to serve as mediator between the model’s internal state and a user’s understanding of it. This is true in cases of local explainability—eg. the importance of a specific feature in a decision tree for a particular prediction—but also when global explainability is the goal—eg. the coefficients of a linear model, along with their meanings and interactions, which provide an overall picture of the model’s behavior. From this perspective, the evaluation of explainability can always be treated as the evaluation of explanations.
3 Intended context
When evaluating any explanation, one critical aspect to be considered is the context in which the explanation is to be used. An explanation that is useful for researchers carefully analyzing and modifying a model’s behavior in a controlled environment may not be useful for a teacher trying to understand in real time why a student is struggling with a particular concept. The intended users for which the explanation has been designed must clearly dictate the criteria used to evaluate it. This is often what is meant by the term “human-centered”, which is used in a commendable effort to distance research from simplistic technocentric approaches, instead emphasizing the importance of people. But identifying “humans” as the target of our XAI efforts is still far too broad.
When considering the requirements that an explanation should aim to fulfill, it is useful to examine both the knowledge and objectives of the intended users [22]. Teachers, for example, may wish to help specific students with the insights gained from an explanation. Their knowledge includes their familiarity with their students and their knowledge of the subject matter. Students, on the other hand, may wish to know why a learning platform is making a specific suggestion in order to gauge its effectiveness. Their knowledge might include their level of familiarity with self-regulation strategies, their current understanding of the subject, and clues from what their peers are doing. Researchers may wish instead to use an explanation to better understand how to improve the model, which can involve reducing bias, improving performance, or identifying and fixing bugs that may be present [27].
Considering users’ knowledge and objectives requires a more nuanced, context-aware approach to evaluation. Some studies have taken a bottom-up approach to understanding these needs. Liao et al. [13] interviewed UX and design practitioners to create an “XAI question bank” with prototypical questions users may wish to have answers to. These include global questions about how a model works, local questions about why a specific prediction was made, counterfactual questions of why not a different prediction, hypothetical questions about how to change the prediction, and more.
A similar approach in education can yield insights into the questions that teachers, students, and other stakeholders may wish to have answered by an explanation. Alternatively, it may be that interactive explanations—perhaps made possible through the abilities of LLMs to answer questions using natural language—will provide different stakeholders with the information that is relevant to them, while also allowing for follow-up questions to better understand explanations [28].
4 Evaluation criteria
The XAI literature has highlighted several criteria to consider when evaluating explanations. Due to the lack of a standardized evaluation framework, these criteria often go by different names, have varying semantic domains, or are haphazardly used interchangeably. Some of the definitions used in the literature implicitly suggest the existence of conceptual dependencies between criteria. However, to the best of our knowledge, they have not been previously described hierarchically. Pulling from both the HCI and ML communities, we here propose a systematic hierarchy of criteria with dependencies between them, as depicted in Figure 1.
Figure 1: Evaluation criteria framework. Edges depict the direction of dependence (A -> B = A depends on B).
The ultimate goal of an explanation is to be useful to the user. In education, the user typically represents a stakeholder in the learning process, such as a teacher, student, parent, or administrator, but it can also be a researcher who develops and improves the model.
Intuitively, in order for an explanation to be useful, it must meet the criterion of intelligibility, which refers to how well it can be understood. This concept has also been called “explicitness” [2] and “comprehensibility” [5]. As discussed in the previous section, the specific context and target user for which an explanation has been developed is crucial to an accurate evaluation of intelligibility. In education, an intelligible explanation is one that can be understood by a student, teacher, or a different stakeholder, depending on its intended context. The term “intelligibility” arose within the HCI community [4] and continues to be the predominant term used by HCI researchers for what the ML community refers to as “explainability” or “interpretability”. We discuss further differences and similarities between these two communities later in this paper.
Just as intuitively—though slightly more contentiously—useful explanations must also be faithful. In this context, faithfulness refers to the level of accuracy with which an explanation reflects the model’s internal state [2, 9]. Faithful explanations have also been called accurate explanations [20] and high-fidelity explanations [5, 17]. Faithful explanations can be thought of as providing a view of the model’s internal causality (what leads to its predictions). In education, a faithful explanation is one that provides accurate insights into why a model has made a specific content prediction, such as a study content recommendation or the detection of learner disengagement.
Unlike with intelligibility, however, there is not universal agreement on the necessity of explanation faithfulness. This disagreement arises from the use of post-hoc explainability techniques—such as LIME [19] and SHAP [16]—that derive explanations from a simplified approximation of a more complex model. Post-hoc explanations don’t directly access a model’s internal causality, but rather provide a justification of predictions after-the-fact. Some argue that a lack of faithfulness can lead to misleading and problematic explanations [20, 24] while others suggest that approximate explanations can be used to achieve “sufficient understanding” for specific users performing specific tasks [14]. We argue that, while perfect faithfulness to a model may not be necessary in all contexts, a high level of faithfulness is nevertheless important for an explanation to be useful.
Note that intelligibility and faithfulness are independent criteria. An explanation can be very intelligible but not particularly faithful, or highly faithful but quite unintelligible. Yet a useful explanation requires both conditions to be present past a minimum threshold.
Another criterion described in the literature is plausibility [9]. A plausible explanation is one that aligns with human intuition. For example, explaining that a model predicts a student is disengaged because they did well on a problem is nonsensical. Cases of an explanation that is faithful but not plausible serve as evidence of a problem with the model itself—perhaps it is overfitted and is picking up on noise in the training data. Because plausibility is important to sensemaking, we argue that it is a prerequisite for intelligibility.
Stability refers to the consistency of an explanation for similar examples [2, 5]. That is, an explanation is stable when it provides similar results for similar inputs. For example, one would expect a learner model to provide similar latent knowledge estimates on a particular knowledge component for students who encountered similar struggles on the same problems. If an explanation is not stable, it is difficult to trust it as a reliable source of information. Stability is also a prerequisite for faithfulness. If an explanation is not adequately stable, it is unlikely to be faithful to the model’s internal state.
5 Bridging perspectives
As noted earlier, our evaluation criteria hierarchy for explanations is informed by two distinct research communities: the ML and HCI communities. While there is much overlap between them, Liao & Varshney [14] have pointed out a tension in the goals and methods used by these two communities. The ML community and the XAI sub-community have primarily focused on technical solutions to the challenge of interpretability, often relying on lower-level explainability methods that suit the needs of engineers. The HCI community, on the other hand, has been more heavily informed by the social and information sciences, which has led to more user-centered approaches often based on participatory design methods.
The terms and definitions used by these communities are illustrative of their differing perspectives. The ML community has settled on terms such as “explainability” and “interpretability”, and has even fostered a growing group of eXplainable AI (XAI) researchers. The coining of the term XAI has been attributed to van Lent et al. [26], who used it to describe a system that can present an “easily understood chain of reasoning” from input, “through the AI’s knowledge and inference”, to the final prediction. The HCI community, on the other hand, prefers the term “intelligibility”, which was originally defined as systems that “represent to their users what they know, how they know it, and what they are doing about it” [4]. Notice the emphasis that explainability places on the prediction process from input to output, contrasted with the pragmatic emphasis on users in the HCI definition. Yet despite these differences, there are more overlaps between these communities than points of divergence.
It may be that at least part of the tension described by Liao & Varshney [14] is the result of a semantic misalignment between the two groups. Technical approaches tend to emphasize explanation faithfulness because they emphasize the role of engineer-researcher as the target user, while socio-behavioral approaches care more about explanation intelligibility that have non-researchers as the end-users. In other words, while both communities are working towards the same goal of making AI understandable by people, they are doing so from different perspectives and with different priorities, which leads them to sometimes talk past each other without realizing it. Vaughan & Wallach [27] argue for a bringing together of these communities to create a more holistic approach to XAI.
It should also be noted that some HCI researchers include aspects of transparency not often considered to be within the realm of explainability as crucial to its goals. These go beyond the internal workings of a model, including explanations of the data used for training, performance metrics, levels of uncertainty, and the types of features it relies on [14, 27]. Among education researchers, Kay et al. [11] make reference to the concept of scrutability in the sense of being able to scrutinize a model or system (with a heavy focus on learners as target users). Full scrutability may require similar aspects of transparency that go beyond the model itself.
6 Evaluation methods
An additional layer above that of which evaluation criteria to use is the question of which methods to use to evaluate explanations. The evaluation method will dictate the specific criteria that can be measured. Using Doshi-Velez & Kim [8] as a guided taxonomy of evaluation methods, we can see how the evaluation criteria framework we have proposed can be used alongside these different methods. Within this taxonomy, the choice of evaluation method depends on the domain-specific needs and the context of intended interpretability.
Doshi-Velez & Kim [8] propose three categories of evaluation methods. In decreasing level of resource complexity, they are:
  • Application-grounded evaluation, which involves human users performing realistic tasks.
  • Human-grounded evaluation, which involves human users performing simplified tasks.
  • Functionally-grounded evaluation, which does not involve humans but rather uses quantifiable properties of explanations as a proxy for interpretability.
An example of application grounded evaluation in education is the way learning dashboards are sometimes evaluated by how well they help instructors understand and provide help to students [21, 25], or work on open learner models (OLMs) that provide students with explanations of a model’s estimates of their understanding [7]. This evaluation method can be used to effectively measure explanation intelligibility (and, by extension, plausibility), but it does not directly tackle the question of explanation faithfulness.
Functionally-grounded evaluation, being the least direct category, makes it difficult to make any claims about either intelligibility or faithfulness. It allows for a proxy measurement of intelligibility by considering properties such as model sparsity or explanation simplicity [17], but it does not capture the specific needs of any end-user. While it can be helpful to consider potential target users while conducting this type of evaluation—ideally realistic stakeholders in education—the results are generally context-agnostic and therefore may lack real-world validity. Stability is perhaps the only criterion that can effectively be evaluated using functionally-grounded evaluation. This method may be most appropriate for preliminary studies in an area without much prior research.
Some forms of human-grounded evaluation, on the other hand, are more likely to capture evidence of explanation faithfulness. Doshi-Velez & Kim [8] identify three examples within this category: binary forced choice, forward simulation, and counterfactual simulation.
In binary forced choice, participants must select which explanation they consider best when presented with multiple options. This method was used in an educational context by Swamy, Du, et al. [23] to gauge which explanations were trusted most by university-level educators. This somewhat approximates a measurement of plausibility by allowing participants to identify explanations that match their intuitions, but it does not truly measure intelligibility. It also does not evaluate faithfulness.
In forward simulation, participants must correctly predict the model’s output given specific inputs. An experiment along these lines was proposed by Baker [3] to test interpretability. This provides a very direct measurement of faithfulness, since an explanation must be faithful in order for the task to be performed accurately. It also serves to measure intelligibility, since participants must understand the explanation to succeed. However, given a sufficiently simple model, it may be possible to succeed at a forward simulation task by only using a model’s parameters as explanation without understanding the purpose, features, or even the domain for which the model was built.
A counterfactual simulation is similar to a forward simulation, but participants must correctly identify how a specific input needs to be changed in order to alter the model’s given output. This also allows for an evaluation of both faithfulness and intelligibility, but the same caveats apply as for forward simulation. The ability to recognize valid counterfactuals has been identified by Cohausz [6] as a key step towards using machine learning to design theoretically sound causal models.
7 Evaluation case study
We now turn to an illustration of the concepts discussed here using an ongoing study of an interpretable neural network for predicting a particular learner behavior. The model in question is a convolutional neural network (CNN), designed to be interpretable via targeted regularization to create binary convolutional filters that more accurately align with the input data [10]. The CNN was trained to predict students’ gaming the system behavior (GTS) on a dataset of interactions with the Cognitive Tutor Algebra system. The details and results of an early version of this model were previously reported in Pinto et al. [18].
7.1 Setting up the questionnaire tasks
To evaluate the level of interpretability of this model, we designed a questionnaire that tasks participants with both a forward simulation and counterfactual simulation. The questionnaire is designed for participants from a wide range of backgrounds—both with and without prior experience in machine learning.
The forward simulation task presents participants with the inputs for a particular instance—that is, the values for each variable for a given “clip” of five consecutive student actions. They must then predict whether the model would label this clip as GTS or not GTS, given the patterns in the convolutional filters.
The counterfactual simulation task again presents participants with the inputs for a specific clip, but this time also providing the model’s predicted label. Participants are asked to identify a single change to the inputs that would alter the model’s prediction. For example, given a series of inputs and the model’s prediction of not GTS, what change to the inputs would result in the model labeling this clip as GTS. Participants select the single correct answer from a series of multiple-choice options.
Figure 2 shows an example from the digital questionnaire platform. The inputs (leftmost blue grid) are presented in a simplified tabular format: a grid with features stacked vertically (labeled v01–v24), with each column representing a separate action in the sequence (labeled 1–5). Blue cells indicate a feature value of 1 (present), while white cells indicate a feature value of 0 (absent). The binary convolutional filters (green grids) are represented in the same manner, but each only depicts three actions to match the kernel size of the model’s convolutional layer.
Figure 2: Example visualization from the questionnaire that serves as the core of the model’s explanation.
This visualization—along with the background information provided—serves as the model’s explanation, showing the patterns that the model has learned to associate with GTS behavior. It presents both a global explanation of the model’s logic (patterns that are indicative of GTS) and explanations of specific outputs (the model’s prediction for a particular clip of student actions). The questionnaire is used as a tool to evaluate the explanations themselves.
For the forward simulation task, the instructions ask the following questions: “would the model identify the following clip of student actions as GTS or not GTS? If GTS, what is the number of the matching model pattern?” In the counterfactual simulation task, we ask “which of the following changes to the input would alter the model’s prediction?” Possible answers for the counterfactual simulation include the addition or removal of specific actions, such as “add v07 to action 2” or “remove v19 at action 4”. For both tasks, we also ask participants to rate their confidence on each question.
7.2 Evaluating the evaluation
The evaluation methods used in this questionnaire clearly fall within the category of human-grounded evaluation in the method taxonomy proposed by Doshi-Velez & Kim [8]—they involve human users performing simplified tasks. As such, they provide measurements of faithfulness and intelligibility. By calculating the average accuracy rate (proportion correct out of total questions) across the entire sample of participants, we can quantify how well the explanations were understood (intelligibility). Because the tasks align so closely with the model’s actions, the accuracy rate also serves to measure how well the explanations reflect the model’s internal state (faithfulness).
However, the caveat provided earlier in regards to forward and counterfactual simulation tasks applies here—participants are not required to understand the specific purpose of the model or the value of its predictions in order to succeed. In fact, while we present an explanation of GTS and the aims of the model as background information, we’ve entirely excluded meaningful feature labels from the explanations. This approach makes it impossible to evaluate explanation plausibility, weakening its claims of evaluating intelligibility beyond a surface-level understanding.
Furthermore, this questionnaire does not claim to target any specific end-users. It has been designed for participants from a wide range of backgrounds, and for no specific purpose other than its completion. We previously highlighted the importance of intended context when evaluating explanations, which is difficult to account for using the simplified tasks of human-grounded methods. An application-grounded evaluation would allow for a better understanding of the specific needs of end-users, but it would also make it difficult to measure faithfulness and would require a more complex and time-consuming study design [8]. When it comes to designing an evaluation, tradeoffs may be necessary.
8 Discussion
Much like the complexity of evaluating the different aspects of a model’s performance, the evaluation of explanations is itself a complex task and cannot be captured in its entirety by any one metric or method. We have aimed to provide an initial framework to guide this daunting but important task. However, much work remains to be done.
We have brought together evaluation criteria described by different communities into a cohesive whole, but they largely remain abstract ideas. In order to be useful in practice, these criteria must be operationalized more concretely in the educational contexts in which we wish to use them. Furthermore, this high-level overview is likely missing key criteria that measure aspects of explanations that are currently not being captured.
For example, when describing the aspects of intelligibility that can be captured by human-grounded evaluation methods, as well as those that may go overlooked by such an approach, we found that we didn’t have the exact language to elucidate our point. It may be that there is an element of intelligibility that requires an additional criterion to fully capture—something along the lines of an explanation’s fidelity to its intended context.
Similarly, the framework’s hierarchical structure itself may benefit from further scrutiny. Edge cases theoretically could exist that don’t perfectly fit, such as the possibility of a highly overfitted model leading to explanations that are faithful but not very stable.
Nevertheless, future research may build on the framework and ideas presented here to create a more comprehensive evaluation framework for explanations. A unified framework should be adaptable to the specific needs of different contexts, should be informed by the perspectives of both the technical ML and human-centered HCI communities, and should be relevant to the needs of stakeholders in education.
9 Conclusion
In this position paper, we have proposed the need for a unified framework for evaluating explanations in the context of XAI. We have reviewed important concepts for better understanding the nature of explanations, including their role as mediators between models and users, the central role played by an explanation’s intended context, and the varied perspectives brought by different research communities. We have further argued that useful explanations require both faithfulness and intelligibility, and have proposed a hierarchy of criteria that brings together concepts previously described in the literature. Finally, we have provided a case study for these criteria using the ongoing evaluation of a neural-network-based learner behavior detector.